OpenStudy (anonymous):

Consider line segment AB whose endpoints are (1, 4) and (4, 8). If this segment is dilated by a scale factor of 0.5 centered at the origin, find the length of the new segment, A’B’.

OpenStudy (anonymous):

5 10 0.65 2.5

OpenStudy (anonymous):

I think we need to use the distance formula for this one. Let me find it

OpenStudy (anonymous):

\[d = \sqrt{(x _{2}-x _{1})^{2}+(y _{1}-y _{2})^{2}}\]

OpenStudy (anonymous):

(1, 4) and (4, 8) So we plug that into the formula, like so: \[d = \sqrt{4-1)^{2}+(8-4)^{2}}\]

OpenStudy (anonymous):

Oh, and it's y2 - y1, not y1 - y2, mistyped it.

OpenStudy (anonymous):

Now we solve what's under the sqrt. 4 - 1 = 3 8 - 4 = 4 sqrt (3)^2 + (4)^2

OpenStudy (anonymous):

sqrt 9 + 16

OpenStudy (anonymous):

sqrt 25

OpenStudy (anonymous):

d = 5

OpenStudy (anonymous):

So now we multiply 5 by 0.5

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

Thank you!!! @jcpd910